# Wolfpack Adversarial Attack for Robust Multi-Agent Reinforcement Learning

### v2

@@ Abstract
Traditional robust methods in multi-agent reinforcement learning (MARL) often struggle against coordinated adversarial attacks in cooperative scenarios. To address this limitation, we propose the Wolfpack Adversarial Attack framework, inspired by wolf hunting strategies, which targets an initial agent and its assisting agents to disrupt cooperation. Additionally, we introduce the Wolfpack-Adversarial Learning for MARL (WALL) framework, which trains robust MARL policies to defend against the proposed Wolfpack attack by fostering system-wide collaboration. Experimental results underscore the devastating impact of the Wolfpack attack and the significant robustness improvements achieved by WALL.
@@ 4.3 Follow-up Agent Group Selection Method
where $\Delta Q_{t}^{tot}\geq 0$ for all $t$ such that $\mathcal{N}_{t,\mathrm{attack}}\neq\emptyset$, because $\tilde{\mathbf{a}}t$ minimizes $Q^{tot}$ for the agent indices selected by $\pi_{\mathrm{adv}}^{\mathrm{WP}}$. Assuming the $i$-th agent is the target of the initial attack, updating $Q^{tot}$ based on $\Delta Q_{t_{\mathrm{init}}}^{tot}$ adjusts each agent’s individual value function $Q^{j}$ to increase $Q^{tot}$ for all $j\neq i\in\mathcal{N}$, in accordance with the credit assignment principle in CTDE algorithms (Sunehag et al., 2017; Rashid et al., 2020), as shown below:
$\tilde{Q}^{i}(\tau_{t_{\mathrm{init}}}^{i},\cdot),=Q^{i}(\tau_{t_{\mathrm{init}}}^{i},\cdot)-\alpha_{\mathrm{lr}}\left.{\partial\Delta Q_{t_{\mathrm{init}}}^{tot}\over\partial Q^{i}(\tau_{t_{\mathrm{init}}}^{i},\mathbf{a})}\right|_{\mathbf{a}=\tilde{\mathbf{a}}_{t_{\mathrm{init}}}},$ (1)
@@ 4.5 WALL: A Robust MARL Algorithm
Similar to other robust MARL methods, we propose the Wolfpack-Adversarial Learning for MARL (WALL) framework, a robust policy designed to counter the Wolfpack attack by performing MARL on the LPA-Dec-POMDP $\tilde{\mathcal{M}}$ with the Wolfpack attacker $\pi_{\mathrm{adv}}^{\mathrm{WP}}$. We utilize well-known CTDE algorithms, including VDN (Sunehag et al., 2017), QMIX (Rashid et al., 2020), and QPLEX (Wang et al., 2020b). Detailed implementations, including loss functions for the planner transformer and the value functions, are provided in Appendix B.2. The proposed WALL framework is illustrated in Fig. 5 and summarized in Algorithm 1.
@@ 5 Experiments
In this section, we evaluate the proposed methods in the Starcraft II Multi-Agent Challenge (SMAC) (Samvelyan et al., 2019) environment, a benchmark in MARL research. Specifically, we compare: (1) the impact of the proposed Wolfpack adversarial attack against other adversarial attacks, and (2) the robustness of the WALL framework in defending against such attacks compared to other robust MARL methods. Also, an ablation study analyzes the effect of the proposed components and hyperparameters on robustness. All results are reported as the mean and standard deviation (shaded areas for graphs and $\pm$ values for tables) across 5 random seeds.
Method Map Name
@@
Natural Vanilla QMIX $98.0\pm 1.52$ $99.2\pm 1.03$ $99.2\pm 1.62$ $97.6\pm 2.10$ $99.2\pm 0.54$ $99.1\pm 1.16$ $98.7\pm 0.56$
RANDOM $\mathbf{99.7\pm 0.56}$ $99.1\pm 1.26$ $99.0\pm 0.88$ $99.2\pm 1.03$ $\mathbf{99.6\pm 0.62}$ $99.3\pm 1.06$ $99.3\pm 0.21$
RARL $97.8\pm 2.09$ $93.8\pm 3.29$ $93.1\pm 17.4$ $95.5\pm 3.75$ $90.6\pm 20.8$ $84.0\pm 33.7$ $92.5\pm 5.39$
RAP $98.8\pm 1.37$ $95.8\pm 4.45$ $99.5\pm 1.08$ $94.7\pm 6.79$ $95.5\pm 12.1$ $84.2\pm 16.9$ $94.7\pm 2.62$
ERNIE $98.2\pm 1.36$ $99.2\pm 1.24$ $99.8\pm 0.44$ $\mathbf{99.8\pm 0.54}$ $98.5\pm 1.78$ $99.2\pm 1.05$ $99.1\pm 0.59$
ROMANCE $96.4\pm 2.97$ $93.6\pm 13.7$ $99.7\pm 0.59$ $99.6\pm 0.62$ $96.4\pm 6.62$ $96.5\pm 4.38$ $97.0\pm 2.11$
WALL (ours) $99.4\pm 0.65$ $\mathbf{99.7\pm 0.82}$ $\mathbf{99.8\pm 0.67}$ $99.3\pm 0.59$ $99.0\pm 2.12$ $\mathbf{99.5\pm 0.62}$ $\mathbf{99.4\pm 0.53}$
Random Attack Vanilla QMIX $80.4\pm 3.24$ $69.6\pm 10.4$ $91.4\pm 4.61$ $69.0\pm 7.19$ $66.4\pm 20.5$ $94.8\pm 3.45$ $78.6\pm 2.22$
RANDOM $90.8\pm 3.87$ $76.4\pm 17.3$ $97.4\pm 0.68$ $80.2\pm 4.92$ $95.4\pm 5.31$ $96.0\pm 2.91$ $89.4\pm 2.54$
RARL $86.8\pm 4.06$ $58.7\pm 15.4$ $89.2\pm 20.4$ $70.2\pm 5.83$ $84.2\pm 5.41$ $79.1\pm 22.2$ $78.0\pm 7.26$
RAP $91.0\pm 4.73$ $69.2\pm 11.1$ $97.8\pm 1.62$ $85.0\pm 11.4$ $86.7\pm 30.3$ $86.6\pm 12.5$ $86\pm 3.74$
ERNIE $83.2\pm 6.94$ $65.2\pm 4.92$ $90.2\pm 9.23$ $76.2\pm 11.8$ $86.0\pm 18.2$ $95.6\pm 3.24$ $82.7\pm 5.27$
ROMANCE $90.2\pm 2.39$ $71.6\pm 10.88$ $99.6\pm 0.68$ $84.8\pm 5.00$ $86.8\pm 16.3$ $94.0\pm 1.96$ $87.8\pm 2.44$
WALL (ours) $\mathbf{94.6\pm 4.53}$ $\mathbf{87.4\pm 1.88}$ $\mathbf{99.8\pm 0.56}$ $\mathbf{95.8\pm 3.45}$ $\mathbf{99.4\pm 1.11}$ $\mathbf{98.6\pm 1.67}$ $\mathbf{95.9\pm 0.54}$
EGA Vanilla QMIX $54.0\pm 7.60$ $66.5\pm 15.5$ $72.4\pm 15.1$ $71.2\pm 20.0$ $70.6\pm 14.8$ $83.0\pm 2.61$ $69.6\pm 3.84$
RANDOM $65.3\pm 3.31$ $70.6\pm 38.6$ $68.8\pm 23.5$ $87.5\pm 5.53$ $84.4\pm 3.12$ $84.5\pm 2.97$ $76.9\pm 5.44$
RARL $62.6\pm 9.40$ $74.4\pm 12.4$ $88.4\pm 17.7$ $78.4\pm 9.16$ $83.4\pm 11.9$ $80.1\pm 11.4$ $77.9\pm 9.21$
RAP $70.4\pm 13.0$ $84.4\pm 7.37$ $83.8\pm 15.8$ $86.2\pm 3.81$ $83.9\pm 16.4$ $80.2\pm 5.41$ $81.5\pm 4.38$
ERNIE $52.4\pm 9.68$ $60.4\pm 20.1$ $83.2\pm 9.75$ $81.6\pm 8.54$ $85.0\pm 4.65$ $93.6\pm 2.26$ $76.0\pm 3.00$
ROMANCE $79.8\pm 2.83$ $85.8\pm 4.60$ $91.0\pm 5.19$ $90.9\pm 4.09$ $87.8\pm 11.7$ $89.6\pm 2.99$ $87.5\pm 1.69$
WALL (ours) $\mathbf{88.6\pm 5.43}$ $\mathbf{87.0\pm 5.45}$ $\mathbf{98.7\pm 0.84}$ $\mathbf{95.8\pm 2.96}$ $\mathbf{94.2\pm 3.85}$ $\mathbf{97.0\pm 1.32}$ $\mathbf{93.6\pm 1.58}$
Wolfpack Adversarial Attack (ours) Vanilla QMIX $39.8\pm 7.52$ $31.0\pm 11.8$ $84.4\pm 4.78$ $11.4\pm 13.3$ $10.4\pm 14.3$ $59.2\pm 4.25$ $39.4\pm 4.57$
RANDOM $60.4\pm 27.4$ $57.4\pm 30.5$ $91.0\pm 3.17$ $40.4\pm 14.8$ $63.6\pm 28.7$ $68.4\pm 19.6$ $63.5\pm 3.14$
RARL $52.4\pm 15.8$ $31.1\pm 20.3$ $90.0\pm 17.4$ $14.2\pm 9.71$ $51.1\pm 36.8$ $75.9\pm 13.6$ $52.5\pm 8.02$
RAP $60.0\pm 10.3$ $37.5\pm 10.4$ $95.4\pm 3.98$ $35.6\pm 14.4$ $47.0\pm 36.9$ $75.7\pm 26.1$ $58.5\pm 5.79$
ERNIE $43.2\pm 13.0$ $35.4\pm 6.25$ $94.8\pm 4.42$ $26.4\pm 10.4$ $26.2\pm 17.3$ $77.0\pm 9.33$ $50.5\pm 6.62$
ROMANCE $62.4\pm 5.16$ $34.8\pm 14.3$ $98.6\pm 0.68$ $28.6\pm 14.2$ $48.8\pm 17.9$ $81.2\pm 4.51$ $59.1\pm 2.08$
WALL (ours) $\mathbf{92.2\pm 3.77}$ $\mathbf{90.8\pm 4.92}$ $\mathbf{99.8\pm 0.56}$ $\mathbf{83.6\pm 5.01}$ $\mathbf{95.0\pm 4.56}$ $\mathbf{98.8\pm 1.62}$ $\mathbf{93.4\pm 1.18}$
@@
Table 1: Average test win rates of robust MARL policies under various attack settings
5.1 Environmental Setup
The SMAC environment serves as a challenging benchmark requiring effective agent cooperation to defeat opponents. We evaluate the proposed method across six scenarios: 2s3z, 3m, 3s_vs_3z, 8m, MMM, and 1c3s5z. We perform parameter searches for the number of follow-up agents $m$, total Wolfpack attacks $K_{\mathrm{WP}}$, and attack duration $t_{\mathrm{WP}}$, using optimal settings for comparisons. To ensure realistic constraints, we set $m$ to $m<\left\lfloor\frac{n-1}{2}\right\rfloor$, where $n$ is the maximum number of allied units. We provide details on environment setups and experimental configurations, including hyperparameter settings, in Appendices A and C. All MARL methods are evaluated on the QMIX baseline, with comparison for other CTDE baselines in Appendix E.1.
@@ 5.1 Environmental Setup
Robust MARL Baselines: To compare the severity of attack baselines and the robustness of policies trained under adversarial attack scenarios, we evaluate QMIX-trained policies under the following attack conditions: Vanilla QMIX, assuming no adversarial attacks; RANDOM, using Random Attack; RARL (Pinto et al., 2017), where adversarial attackers tailor attacks to the learned policy; RAP (Vinitsky et al., 2020), an extension of RARL that uniformly samples attackers to prevent overfitting and introduce diversity; ROMANCE (Yuan et al., 2023), an RAP extension countering diverse EGA attacks; ERNIE (Bukharin et al., 2024), enhancing robustness via adversarial regularization in observations and actions; and the proposed WALL.
All robust MARL methods follow author-provided methodologies and parameters. Further details on the MARL baselines are available in Appendix D. All policies are trained for 3M timesteps, starting from a pretrained Vanilla QMIX model trained for 1M timesteps.
@@
5.2 Performance Comparison

### v3

@@ Abstract
Traditional robust methods in multi-agent reinforcement learning (MARL) often struggle against coordinated adversarial attacks in cooperative scenarios. To address this limitation, we propose the Wolfpack Adversarial Attack framework, inspired by wolf hunting strategies, which targets an initial agent and its assisting agents to disrupt cooperation. Additionally, we introduce the Wolfpack-Adversarial Learning for MARL (WALL) framework, which trains robust MARL policies to defend against the proposed Wolfpack attack by fostering systemwide collaboration. Experimental results underscore the devastating impact of the Wolfpack attack and the significant robustness improvements achieved by WALL. Our code is available at https://github.com/sunwoolee0504/WALL.
@@ 4.3 Follow-up Agent Group Selection Method
where $\Delta Q_{t}^{tot}\geq 0$ for all $t$ such that $\mathcal{N}_{t,\mathrm{attack}}\neq\emptyset$, because $\tilde{\mathbf{a}}_{t}$ minimizes $Q^{tot}$ for the agent indices selected by $\pi_{\mathrm{adv}}^{\mathrm{WP}}$. Assuming the $i$-th agent is the target of the initial attack, updating $Q^{tot}$ based on $\Delta Q_{t_{\mathrm{init}}}^{tot}$ adjusts each agent’s individual value function $Q^{j}$ to increase $Q^{tot}$ for all $j\neq i\in\mathcal{N}$, in accordance with the credit assignment principle in CTDE algorithms (Sunehag et al., 2017; Rashid et al., 2020), as shown below:
$\tilde{Q}^{i}(\tau_{t_{\mathrm{init}}}^{i},\cdot)=Q^{i}(\tau_{t_{\mathrm{init}}}^{i},\cdot)-\alpha_{\mathrm{lr}}\left.{\partial\Delta Q_{t_{\mathrm{init}}}^{tot}\over\partial Q^{i}(\tau_{t_{\mathrm{init}}}^{i},\mathbf{a})}\right|_{\mathbf{a}=\tilde{\mathbf{a}}_{t_{\mathrm{init}}}},$ (1)
@@ 4.5 WALL: A Robust MARL Algorithm
Similar to other robust MARL methods, we propose the Wolfpack-Adversarial Learning for MARL (WALL) framework, a robust policy designed to counter the Wolfpack attack by performing MARL on the LPA-Dec-POMDP $\tilde{\mathcal{M}}$ with the Wolfpack attacker $\pi_{\mathrm{adv}}^{\mathrm{WP}}$. While the proposed Wolfpack framework is broadly applicable to most CTDE algorithms, we primarily applied it to well-known value-based CTDE methods, including QMIX (Rashid et al., 2020), VDN (Sunehag et al., 2017), and QPLEX (Wang et al., 2020b). Detailed implementations, including loss functions for the planner Transformer and the value functions, are provided in Appendix B.2. The proposed WALL framework is illustrated in Fig. 5 and summarized in Algorithm 1.
@@
Figure 6: MARL benchmarks used in our experiments: (a) PP_3/1 and (b) PP_6/2 in MPE, and (c) 8m and (d) MMM scenarios in SMAC.
@@ 5 Experiments
In this section, we evaluate the proposed methods on two standard benchmarks in MARL research: the Multi-Agent Particle Environment (MPE) (Lowe et al., 2017) and the StarCraft II Multi-Agent Challenge (SMAC) (Samvelyan et al., 2019), as illustrated in Fig. 6. Specifically, we compare: (1) the impact of the proposed Wolfpack adversarial attack against other adversarial attacks, and (2) the robustness of the WALL framework in defending against such attacks compared to other robust MARL methods. Also, an ablation study analyzes the effect of the proposed components and hyperparameters on robustness. All results are reported as the mean and standard deviation (shaded areas for graphs and $\pm$ values for tables) across 5 random seeds. Our code is available at https://github.com/sunwoolee0504/WALL.
5.1 Environmental Setup
The MPE environment provides a multi-agent setting where agents interact through simple physical dynamics. We conduct experiments on three predator-prey (PP) scenarios with varying agent-to-target ratios: PP_3/1, PP_6/2, and PP_9/3. In these tasks, multiple predator agents must coordinate to capture one or more prey agents moving adversarially. The SMAC environment serves as a challenging benchmark requiring effective agent cooperation to defeat opponents. We evaluate the proposed method across six scenarios: 2s3z, 3m, 3s_vs_3z, 8m, MMM, and 1c3s5z. We perform parameter searches for the number of follow-up agents $m$, total Wolfpack attacks $K_{\mathrm{WP}}$, and attack duration $t_{\mathrm{WP}}$, using optimal settings for comparisons. To ensure realistic constraints, we set $m$ to $m<\left\lfloor\frac{n-1}{2}\right\rfloor$, where $n$ is the maximum number of allied units. We provide details on environment setups and experimental configurations, including hyperparameter settings, in Appendices A and C. All MARL methods are evaluated on the QMIX baseline, with comparison for other CTDE baselines in Appendix E.1.
Method Scenario
PP_3/1
PP_6/2
PP_9/3
Mean
Natural Vanilla QMIX $165.4\pm 1.3$ $538.8\pm 5.5$ $661.9\pm 5.4$ $455.4\pm 2.1$
RANDOM $178.3\pm 1.0$ $663.8\pm 4.0$ $666.9\pm 3.5$ $503.0\pm 1.3$
ROMANCE $175.0\pm 0.7$ $648.6\pm 3.9$ $721.4\pm 7.5$ $515.0\pm 1.4$
WALL (ours) $\mathbf{202.9\pm 1.7}$ $\mathbf{675.0\pm 3.8}$ $\mathbf{802.5\pm 5.7}$ $\mathbf{560.1\pm 1.7}$
Radom Attack Vanilla QMIX $158.9\pm 2.1$ $522.6\pm 2.4$ $656.4\pm 3.0$ $445.9\pm 2.0$
RANDOM $170.1\pm 1.3$ $638.7\pm 2.0$ $657.9\pm 2.2$ $488.9\pm 1.9$
ROMANCE $173.4\pm 1.4$ $624.2\pm 3.5$ $701.7\pm 4.0$ $499.7\pm 2.4$
WALL (ours) $\mathbf{202.9\pm 1.7}$ $\mathbf{654.0\pm 2.1}$ $\mathbf{802.5\pm 2.1}$ $\mathbf{553.1\pm 1.7}$
EGA Vanilla QMIX $153.7\pm 2.0$ $499.8\pm 0.6$ $585.1\pm 1.7$ $412.8\pm 1.4$
RANDOM $157.6\pm 1.9$ $604.6\pm 3.2$ $589.7\pm 1.7$ $450.6\pm 1.4$
ROMANCE $167.1\pm 2.4$ $605.1\pm 2.7$ $594.4\pm 1.3$ $455.5\pm 1.5$
WALL (ours) $\mathbf{166.5\pm 1.7}$ $\mathbf{685.5\pm 5.0}$ $\mathbf{682.7\pm 0.8}$ $\mathbf{511.5\pm 1.4}$
Wolfpack Adversarial Attack (ours) Vanilla QMIX $135.7\pm 1.5$ $415.6\pm 8.4$ $551.7\pm 7.0$ $367.6\pm 5.0$
RANDOM $157.5\pm 1.4$ $571.0\pm 9.0$ $624.3\pm 7.7$ $450.9\pm 8.4$
ROMANCE $159.3\pm 1.5$ $554.6\pm 9.9$ $570.3\pm 7.5$ $428.0\pm 9.4$
WALL (ours) $\mathbf{171.5\pm 1.5}$ $\mathbf{599.0\pm 9.4}$ $\mathbf{698.1\pm 9.1}$ $\mathbf{489.5\pm 0.6}$
Table 1: Average cumulative rewards of robust MARL policies under various attack settings in the MPE environments.
Method Scenario
@@ 5.1 Environmental Setup
Natural Vanilla QMIX $98.0\pm 1.5$ $99.2\pm 1.0$ $99.2\pm 1.6$ $97.6\pm 2.1$ $99.2\pm 0.5$ $99.1\pm 1.1$ $98.7\pm 0.5$
RANDOM $\mathbf{99.7\pm 0.5}$ $99.1\pm 1.2$ $99.0\pm 0.8$ $99.2\pm 1.0$ $\mathbf{99.6\pm 0.6}$ $99.3\pm 1.0$ $99.3\pm 0.2$
RARL $97.8\pm 2.0$ $93.8\pm 3.2$ $93.1\pm 17.4$ $95.5\pm 3.7$ $90.6\pm 20.8$ $84.0\pm 33.7$ $92.5\pm 5.3$
RAP $98.8\pm 1.3$ $95.8\pm 4.4$ $99.5\pm 1.0$ $94.7\pm 6.7$ $95.5\pm 12.1$ $84.2\pm 16.9$ $94.7\pm 2.6$
ERNIE $98.2\pm 1.3$ $99.2\pm 1.2$ $99.8\pm 0.4$ $\mathbf{99.8\pm 0.5}$ $98.5\pm 1.7$ $99.2\pm 1.0$ $99.1\pm 0.5$
ROMANCE $96.4\pm 2.9$ $93.6\pm 13.7$ $99.7\pm 0.5$ $99.6\pm 0.6$ $96.4\pm 6.6$ $96.5\pm 4.3$ $97.0\pm 2.1$
WALL (ours) $99.4\pm 0.6$ $\mathbf{99.7\pm 0.8}$ $\mathbf{99.8\pm 0.7}$ $99.3\pm 0.6$ $99.0\pm 2.1$ $\mathbf{99.5\pm 0.6}$ $\mathbf{99.4\pm 0.5}$
Random Attack Vanilla QMIX $80.4\pm 3.2$ $69.6\pm 10.4$ $91.4\pm 4.6$ $69.0\pm 7.1$ $66.4\pm 20.5$ $94.8\pm 3.4$ $78.6\pm 2.2$
RANDOM $90.8\pm 3.8$ $76.4\pm 17.3$ $97.4\pm 0.6$ $80.2\pm 4.9$ $95.4\pm 5.3$ $96.0\pm 2.9$ $89.4\pm 2.5$
RARL $86.8\pm 4.0$ $58.7\pm 15.4$ $89.2\pm 20.4$ $70.2\pm 5.8$ $84.2\pm 5.4$ $79.1\pm 22.2$ $78.0\pm 7.2$
RAP $91.0\pm 4.7$ $69.2\pm 11.1$ $97.8\pm 1.6$ $85.0\pm 11.4$ $86.7\pm 30.3$ $86.6\pm 12.5$ $83.0\pm 3.7$
ERNIE $83.2\pm 6.9$ $65.2\pm 4.9$ $90.2\pm 9.2$ $76.2\pm 11.8$ $86.0\pm 18.2$ $95.6\pm 3.2$ $82.7\pm 5.2$
ROMANCE $90.2\pm 2.3$ $71.6\pm 10.8$ $99.6\pm 0.6$ $84.8\pm 5.0$ $86.8\pm 16.3$ $94.0\pm 1.9$ $87.8\pm 2.4$
WALL (ours) $\mathbf{94.6\pm 4.5}$ $\mathbf{87.4\pm 1.8}$ $\mathbf{99.8\pm 0.5}$ $\mathbf{95.8\pm 3.4}$ $\mathbf{99.4\pm 1.1}$ $\mathbf{98.6\pm 1.6}$ $\mathbf{95.9\pm 0.5}$
EGA Vanilla QMIX $54.0\pm 7.6$ $66.5\pm 15.5$ $72.4\pm 15.1$ $71.2\pm 20.0$ $70.6\pm 14$ $83.0\pm 2.6$ $69.6\pm 3.8$
RANDOM $65.3\pm 3.3$ $70.6\pm 38.6$ $68.8\pm 23.5$ $87.5\pm 5.5$ $84.4\pm 3.1$ $84.5\pm 2.9$ $76.9\pm 5.4$
RARL $62.6\pm 9.4$ $74.4\pm 12.4$ $88.4\pm 17.7$ $78.4\pm 9.1$ $83.4\pm 11.9$ $80.1\pm 11.4$ $77.9\pm 10.0$
RAP $70.4\pm 13.0$ $84.4\pm 7.3$ $83.8\pm 15.8$ $86.2\pm 3.8$ $83.9\pm 16.4$ $80.2\pm 5.4$ $81.5\pm 4.3$
ERNIE $52.4\pm 9.6$ $60.4\pm 20.1$ $83.2\pm 9.7$ $81.6\pm 8.5$ $85.0\pm 4.6$ $93.6\pm 2.2$ $76.0\pm 3.0$
ROMANCE $79.8\pm 2.8$ $85.8\pm 4.6$ $91.0\pm 5.1$ $90.9\pm 4.0$ $87.8\pm 11.7$ $89.6\pm 2.9$ $87.5\pm 1.6$
WALL (ours) $\mathbf{88.6\pm 5.4}$ $\mathbf{87.0\pm 5.4}$ $\mathbf{98.7\pm 0.8}$ $\mathbf{95.8\pm 2.9}$ $\mathbf{94.2\pm 3.8}$ $\mathbf{97.0\pm 1.3}$ $\mathbf{93.6\pm 1.5}$
Wolfpack Adversarial Attack (ours) Vanilla QMIX $39.8\pm 7.5$ $31.0\pm 11.8$ $84.4\pm 4.7$ $11.4\pm 13.3$ $10.4\pm 14.3$ $59.2\pm 4.2$ $39.4\pm 4.5$
RANDOM $60.4\pm 27.4$ $57.4\pm 30.5$ $91.0\pm 3.1$ $40.4\pm 14.8$ $63.6\pm 28.7$ $68.4\pm 19.6$ $63.5\pm 3.1$
RARL $52.4\pm 15.8$ $31.1\pm 20.3$ $90.0\pm 17.4$ $14.2\pm 9.7$ $51.1\pm 36.8$ $75.9\pm 13.6$ $52.5\pm 8.0$
RAP $60.0\pm 10.3$ $37.5\pm 10.4$ $95.4\pm 3.9$ $35.6\pm 14.4$ $47.0\pm 36.9$ $75.7\pm 26.1$ $58.5\pm 5.7$
ERNIE $43.2\pm 13.0$ $35.4\pm 6.2$ $94.8\pm 4.4$ $26.4\pm 10.4$ $26.2\pm 17.3$ $77.0\pm 9.3$ $50.5\pm 6.6$
ROMANCE $62.4\pm 5.1$ $34.8\pm 14.3$ $98.6\pm 0.6$ $28.6\pm 14.2$ $48.8\pm 17.9$ $81.2\pm 4.5$ $59.1\pm 2.0$
WALL (ours) $\mathbf{92.2\pm 3.7}$ $\mathbf{90.8\pm 4.9}$ $\mathbf{99.8\pm 0.5}$ $\mathbf{83.6\pm 5.0}$ $\mathbf{95.0\pm 4.5}$ $\mathbf{98.8\pm 1.6}$ $\mathbf{93.4\pm 1.1}$
@@
Table 2: Average test win rates of robust MARL policies under various attack settings in the SMAC environments.
@@
Robust MARL Baselines: To compare the severity of attack baselines and the robustness of policies trained under adversarial attack scenarios, we evaluate QMIX-trained policies under the following attack conditions: Vanilla QMIX, assuming no adversarial attacks; RANDOM, using Random Attack; RARL (Pinto et al., 2017), where adversarial attackers tailor attacks to the learned policy; RAP (Vinitsky et al., 2020), an extension of RARL that uniformly samples attackers to prevent overfitting and introduce diversity; ROMANCE (Yuan et al., 2023), an RAP extension countering diverse EGA attacks; ERNIE (Bukharin et al., 2024), enhancing robustness via adversarial regularization in observations and actions; and the proposed WALL. All robust MARL methods follow author-provided methodologies and parameters. Further details on the MARL baselines are available in Appendix D. All policies are trained for 3M timesteps, starting from a pretrained Vanilla QMIX model trained for 1M timesteps.
@@
5.2 Performance Comparison in MPE and SMAC
Table 1 presents the average cumulative rewards over the last 100 episodes under various attack settings in the MPE environments. The results show that the proposed Wolfpack adversarial attack is significantly more disruptive than existing methods such as EGA and Random Attack. For example, for Vanilla QMIX, the average cumulative reward across the three predator-prey scenarios drops by $455.4-367.6=87.8$ under the proposed Wolfpack attack, compared to $455.4-412.8=42.6$ under EGA and $455.4-445.9=9.5$ under Random Attack. These results demonstrate that the Wolfpack attack imposes a much more severe degradation in policy performance. In contrast, the proposed WALL framework consistently achieves the best performance across all attack types. Notably, in the MPE scenarios, WALL outperforms all baselines not only under adversarial attacks but also in the natural setting, demonstrating superior policy quality even without external threats.